# Decoding Open-Ended Information Seeking Goals from Eye Movements in Reading

### v1

@@ 5.1 Data Splits
Following prior work on predictive modeling with OneStop (e.g. Shubi et al., 2024b), we use a 10-fold cross-validation procedure where each of the 10 data splits has a training, validation, and a test set. The validation set and the test set are each partitioned into three disjoint parts that capture three different levels of model generalization, ordered by task difficulty:
  • New Participant: During training, eye-tracking data is available for the paragraph but not for the participant.
  • New Item: Training eye-tracking data is available for the participant but not for the paragraph.
  • New Item & Participant: Neither the participant nor the paragraph was in the training data.
Each data split includes 64% of the trials in the training set, 17% in the validation set, and 19% in the test set – divided into 9% New Participant, 9% New Item, and 1% New Item & Participant. When aggregated across the 10 splits, 90% of the trials appear in both New Participant and New Item evaluation regimes, and 10% appear in the New Item & Participant regime.
@@
We ensure that in each split into training, validation, and test sets, there is a balanced distribution of question types, i.e. $q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}$ for each text in each part of the split. Hyperparameter optimization and model selection are conducted separately for each split. We assume that the evaluation regime with respect to novelty of items and participants at test time is unknown, and therefore model selection is performed using the entire validation set within each split. Further details on the hyperparameter search space, training, and evaluation procedures are provided in Appendix A.
Type Model All Different Spans Same Span
Baselines Chance / Majority $33.0\pm 0.4_{\text{+++}}$ $55.3\pm 0.4_{\text{+++}}$ $49.9\pm 0.4_{\text{+++}}$
@@
Discriminative Haller RNN (Haller et al., 2022) $41.8\pm 0.6^{***}_{\text{+++}}$ $65.6\pm 0.5^{***}_{\text{+++}}$ $52.1\pm 0.5^{**}_{\text{+++}}$
RoBERTEye-Fixations (Shubi et al., 2024b) $\mathbf{49.3}\pm 0.3^{***}$ $\mathbf{70.9}\pm 0.5^{***}$ $\mathbf{57.3}\pm 0.5^{***}$
@@
DalEye-Llama $38.1\pm 0.2^{***}_{\text{+++}}$ $60.3\pm 0.7^{***}_{\text{+++}}$ $50.9\pm 0.2_{\text{+++}}$
Table 1: Question classification test accuracy aggregated across 10 cross-validation splits, with 95% confidence intervals. ‘All’ are the results for three-way classification, ‘Different Spans’ for distinguishing questions with different critical spans, and ‘Same Span’ for distinguishing questions with the same critical span, where the critical span is the information in the passage that is necessary for answering the question. A majority question choice is identical to chance because the data is balanced across questions. As the data is not i.i.d (each participant reads multiple paragraphs, and each paragraph is read by multiple participants) we test for differences between models using a linear mixed effects model with random intercepts and slopes for participants and paragraphs: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$ with the Julia MixedModels package (Bates et al., 2022). The accuracy of all the models is compared to chance performance, where significant gains over this baseline are marked with ‘*’ $p<0.05$, ‘**’ $p<0.01$ and ‘***’ $p<0.001$ in superscript. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked with ‘+’ $p<0.05$, ‘++’ $p<0.01$ and ‘$+$’ $p<0.001$ in subscript.
5.2 Evaluation
5.2.1 Question Classification
We evaluate all the models on the question classification task and report classification accuracy. As described in Section 4, the discriminative models assign a probability to a triplet $\langle P,q^{P},E^{P}_{S}\rangle$, and classification is performed by selecting the candidate question with the highest probability. For the generative models, we select the question with the highest model assigned log-likelihood.
In addition to the overall three-way classification accuracy, we also report a breakdown of the model predictions into two fine-grained evaluations:
Different critical spans
This evaluation tests the model’s ability to distinguish between questions with different critical spans, reducing the classification to a binary decision between $q^{P}_{1,c_{1}}$ and $\{q^{P}_{2,c_{2}},q^{P}_{3,c_{2}}\}$. Note that random choice accuracy in this evaluation is approximately 55%, rather than 50%, due to a Monty Hall-type advantage arising from grouping two questions under one critical span (see visualization in Figure 4 of Appendix B).
Same critical span
This is a more challenging evaluation for differentiating between the two questions for $c_{2}$: $q^{P}_{2,c_{2}}$ versus $q^{P}_{3,c_{2}}$. Note that we disregard the question whose critical span is $c_{1}$. The corresponding chance accuracy is 50%.
5.2.2 Question Reconstruction
The generative models are further evaluated for the quality of the reconstructed questions. As with other generation tasks, evaluating reconstruction quality is a major challenge. Here, we use several heuristic evaluations for both the surface form and semantic content of the generated questions.
The question word
We check if the generated question word matches the question word of the ground truth question. The possible question words are What, When, Where, Who, Whom, Which, Whose, Why, and How. An additional category is Other, reserved for all other words starting a question (7% of all the questions in the data). The agreement with the ground truth questions is measured using pairwise Cohen’s Kappa (Cohen, 1960).
BLEU
The BLEU surface similarity score Papineni et al. (2002) between the generated question and the ground truth question.
UIUC question category
We check if the generated question is of the same semantic category as the ground truth question. To this end, we use the question-type taxonomy suggested by Li and Roth (2002). We use the main categories of the taxonomy: entities, humans, numeric and locations, and the subcategories of the category ‘descriptions’: manner, reasoning, definition and description due to their high frequency in the data. We use GPT-4o (Hurst et al., 2024) to categorize the questions. Similarly to question words, we measure agreement in category assignments using pairwise Cohen’s Kappa. Examples and further details on the classification process, including agreement with human annotations, are provided in Appendix D.
BertScore
(Zhang et al., 2020) provides a cosine similarity score between the generated and ground truth questions based on contextual embeddings extracted from RoBERTa.
QA accuracy
We further propose a downstream evaluation method with the following logic: the closer the generated question is to the true question, the easier it should be to select the correct answer for the true question when replaced with the generated question. We implement this evaluation using a RoBERTa Large model fine-tuned for multiple choice QA on the RACE reading comprehension dataset (Lai et al., 2017). This model was shown to have a high accuracy of 86% on multiple choice QA on OneStopQA (Malmaud et al., 2020). We chose this model because OneStopQA is not in its training data, which is likely the case for newer models, thus avoiding possible data contamination. In this evaluation, we take into account only the questions that RoBERTa answers correctly given the ground truth question. We consider a question to be a valid reconstruction of the ground truth question, if and only if given this question, the paragraph, and the four answers for the ground truth question, the model selects the correct answer.
All five methods provide scores for the model generated questions. To facilitate the interpretation of the quality of these scores, we benchmark them against the following baseline questions.
Arbitrary questions
We generate two arbitrary questions for each paragraph. The first is a question generated by GPT-4o, using the prompt in Section E.7. The second is a question generated by a finetuned Llama 3.1 model trained to decode the correct question without eye movement data.
Incorrect questions
The two additional human-composed questions for each paragraph available in OneStopQA. We provide separate evaluations for the same critical span and the different critical span cases relative to the ground truth question.
6 Results
Figure 3: Question reconstruction evaluations of the DalEye-Llama model for (1) the identity of the generated question word, (2) BLEU score (3) the UIUC semantic category of the question, (4) BERTScore, and (5) downstream QA accuracy based on the answer selection of a multiple-choice question answering model. DalEye-Llama performance is benchmarked against two types of human-composed questions (for a different critical span and for the same span) and against arbitrary questions generated with GPT-4o and Llama 3.1. Presented are means with 95% confidence intervals (bootstrapped, $n=1000$) for three generalization levels to new participants and new readers.
6.1 Question Classification
Table 1 presents the aggregated test accuracy across the 10 data splits for the three-way ‘All’ classification, as well as a breakdown into questions with Different Spans, and questions with the Same Span evaluations. A division of the results into the New Item, New Participant, and New Item & Participant regimes are reported in Tables 2, 3 and 4 in Appendix C. Validation set results are reported in Tables 5, 6, 7 and 8 in Appendix C.
Baselines Both baseline models perform at chance level across all three evaluations. The reason for this failure likely lies in the inherent noise in the distribution of reading times over the paragraph, which deems heuristic methods that rely on this distribution to be ineffective.
@@ 6.1 Question Classification
Models The discriminative models perform above chance, and outperform the generative models on all the evaluations. This outcome is not surprising, as differently from the discriminative models, the generative models were not explicitly trained for the question classification task. RoBERTEye-Fixations achieves the highest accuracy across all three evaluation regimes. Additionally, in Table 2 of Appendix C we observe that this performance advantage is consistent across the New Item, New Participant and New Item & Participant evaluations. Notably, it is also the only model which is substantially above chance in the Same Span regime with $57.3\%$ accuracy. This confirms that distinguishing between questions that ask about the same portion of the paragraph is indeed much more challenging than distinguishing between questions over different parts of the passage. At the same time, the modest success of RoBERTEye on the Same Span evaluation speaks to the fine-grained nature of the information that can be extracted in some cases about the reader’s information seeking goals. In the generative models category, DalEye-Llama outperforms DalEye-LLaVA, where the latter is at chance level.
@@ 6.2 Question Reconstruction
As the DalEye-LLaVA classification results are close to chance (Table 1), we focus on evaluating the question reconstruction quality of the DalEye-Llama model. Figure 3 presents the evaluations, which exhibit largely consistent patterns across the different metrics. Table 9 in Appendix D presents these results in numerical form. Importantly, in the New Participant evaluation, DalEye-Llama generated questions outperform both Arbitrary GPT-4o and Llama generated questions as well as Incorrect Human Questions in both question categories on all the five metrics. This suggests that DalEye-Llama generalizes relatively well to new participants when the textual item appeared in the training set.
However, DalEye-Llama performance drops both in absolute terms and relative to the baselines in the more challenging regimes with New Items. Notably, the DalEye-Llama and the arbitrarily model-generated questions tend to perform worse than the incorrect human questions, which exhibit the highest similarity with the ground truth questions. Specifically, the incorrect human question - same critical span, has the highest BLEU, BERTScore, and QA accuracy, which is likely the result of sharing the critical span textual content with the ground truth question. The incorrect human question - different critical span, has the highest question word and question category similarities. This may be related to a tendency of different questions over the same span not to have the exact same question word and type, whereas different human questions over different spans do not constrain each other in these regards. Figure 10 in section E.6 presents an example with a paragraph, the ground truth question, four DalEye-Llama generated questions in the New Item regime and four in the New Participant regime, each leading to a different answer choice by the QA model. We leave the generalization improvement of DalEye-Llama and DalEye-LLaVA as avenues for future work.
@@ 7 Related Work
We build on three lines of work in NLP and psycholinguistics: (i) multimodal language models; (ii) goal driven reading; and (iii) predictive models of cognitive state from eye movements in reading.
Multimodal Language Models
Multimodal language models extend the capabilities of LLMs beyond text by incorporating additional input modalities (Wu et al., 2023) including visual data (Li et al., 2019, 2024a, among others), time-series signals (Gruver et al., 2023; Liang et al., 2024; Jin et al., 2023; Kawarada et al., 2024), and physiological and cognitive signals—such as EEG and ECG (Hollenstein et al., 2021a; Li et al., 2024b; Mishra et al., 2024; Yang et al., 2025).
Eye movements have been integrated into language models for various tasks (Tomanek et al., 2010b; Tokunaga et al., 2017a; Hollenstein et al., 2022; Shubi et al., 2024b, among others). These approaches typically use encoder-based models, leveraging features derived from the eye movement scanpath. Here, we depart from existing work by integrating eye movements in reading into generative, decoder-based language models, exploring both text- and embedding-based methods for conditioning text generation on eye movements.
Analyses of Goal Based Reading
@@
While most prior work in the psychology of reading focuses on ordinary reading for general comprehension Rayner et al. (2012), goal or task based reading has been acknowledged as a key frontier in this area (Radach and Kennedy, 2004). Several studies found differences in eye movements between ordinary reading and tasks that define a reading manner: skimming, speed reading and proofreading (Just et al., 1982; Kaakinen and Hyönä, 2010; Schotter et al., 2014; Strukelj and Niehorster, 2018; Chen et al., 2023). Additional studies have found differences between ordinary reading and tasks of target word search (Rayner and Raney, 1996; Rayner and Fischer, 1996), word verification (Radach et al., 2008), and topic search (White et al., 2015). Prior work also analyzed eye movements during human linguistic tasks, such as named entity annotation (Tomanek et al., 2010a; Tokunaga et al., 2017b). Differences in reading patterns were also found when readers adopt different perspectives on a given text (Kaakinen et al., 2002; Kaakinen and Hyönä, 2008), or given different learning goals (Rothkopf and Billington, 1979). Overall, these works consistently found a systematic influence of the reading task on the resulting reading patterns.
@@
Our work builds most directly on analyses in Kaakinen et al. (2015), Malmaud et al. (2020), Hahn and Keller (2023) and Shubi and Berzak (2023) who compared eye movements in ordinary reading and open-ended information seeking. These studies showed task conditioning effects on reading patterns, especially with respect to task-critical information in the text. Similarly to these studies, we formulate the information seeking task as a reading comprehension question specific to the text. We further leverage the previously observed task conditioning effects for automatic discrimination between different tasks over the same text.
Decoding Cognitive State
With the rise of modern machine learning and NLP, several studies took advantage of eye movements in reading for prediction tasks about language-related aspects of the reader and their cognitive state. These include studies on prediction of the reader’s linguistic background (Berzak et al., 2017; Reich et al., 2022; Skerath et al., 2023), language proficiency (Berzak et al., 2018), reading comprehension (Ahn et al., 2020; Reich et al., 2022; Mézière et al., 2023; Shubi et al., 2024b, among others), and subjective readability (Reich et al., 2022).

### v2

@@ 5.1 Data Splits
The OneStop eye tracking data is not i.i.d., each participant reads multiple paragraphs, and each paragraph is read by multiple participants. Following prior work on predictive modeling with OneStop (e.g. Shubi et al., 2024), we use a 10-fold cross-validation procedure where each of the 10 data splits has a training, validation, and a test set. To account for the structured nature of the data, the validation set and the test set are each partitioned into three disjoint parts that capture three different levels of model generalization, ordered by task difficulty:
New Participant: Training eye tracking data is available for the text but not for the participant.
New Text: Training eye tracking data is available for the participant but not for the text.
New Text & Participant: Neither the participant nor the text was in the training data.
@@
Each data split includes 64% of the trials in the training set, 17% in the validation set, and 19% in the test set – divided into 9% New Participant, 9% New Text, and 1% New Text & Participant. When aggregated across the 10 splits, 90% of the trials appear in both New Participant and New Text evaluation regimes, and 10% appear in the New Text & Participant regime. We ensure that in each split into training, validation, and test sets, there is a balanced distribution of question types, i.e. $q_{1,c_{1}},q_{2,c_{2}},q_{3,c_{2}}$ for each text in each part of the split. Hyperparameter optimization and model selection are conducted separately for each split. We assume that the evaluation regime with respect to novelty of items and participants at test time is unknown, and therefore model selection is performed using the entire validation set within each split. Further details on the hyperparameter search space, training, and evaluation procedures are provided in Appendix˜C.
5.2 Question Selection
We evaluate all the models on the question selection task and report selection accuracy. As described in Section˜4, the discriminative models assign a probability to a triplet $\langle T,q\in Q_{T},E_{P,T}\rangle$, and select the candidate question with the highest probability. For the generative models, we select the question with the highest model-assigned log-likelihood. In addition to an All overall three-way selection accuracy, we also report a breakdown of the model predictions into two fine-grained evaluations:
Different critical spans This evaluation tests the model’s ability to distinguish between questions with different critical spans, reducing the selection problem to a binary decision between $q_{1,c_{1}}$ and $\{q_{2,c_{2}},q_{3,c_{2}}\}$. Note that random choice accuracy in this evaluation is approximately 55%, rather than 50%, due to a Monty Hall-type setup arising from grouping two questions under one critical span (see Table˜1).
Same critical span This is a more challenging evaluation for differentiating between the two questions for $c_{2}$: $q_{2,c_{2}}$ versus $q_{3,c_{2}}$. Note that here we disregard the question whose critical span is $c_{1}$. The corresponding chance accuracy is 50%.
Table 1: Overview of (a) the three-way question selection task and the breakdown of this task into (b) questions with different critical spans and (c) questions with the same critical span. Chance accuracy under the three evaluation regimes: (a) ${\color[rgb]{0.265625,0.59375,0.1640625}3}/9$ (33.3%), (b) ${\color[rgb]{0.265625,0.59375,0.1640625}5}/9$ (55.5%), (c) ${\color[rgb]{0.265625,0.59375,0.1640625}2}/4$ (50%).
(a) All (b) Different critical spans (c) Same critical span
Predicted $\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$} $\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$} $\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$} $\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$}
True
$\color[rgb]{0.9765625,0.56640625,0.40625}q_{1,c_{1}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{2,c_{2}}$}
$\color[rgb]{0.08203125,0.49609375,0.49609375}q_{3,c_{2}}$}
5.3 Question Reconstruction
The generative models are further evaluated for the quality of the reconstructed questions. As with other generation tasks, evaluating reconstruction quality is a major challenge. Here, we use several heuristic evaluations for both the surface form and the semantic content of the generated questions.
The question word We check if the generated question word matches the question word of the ground truth question. The possible question words are What, When, Where, Who, Whom, Which, Whose, Why, and How. An additional category is Other, reserved for all other words starting a question (7% of the questions). The agreement with the ground truth questions is measured using pairwise Cohen’s Kappa (Cohen, 1960).
UIUC question category We test if the generated question has the same semantic category as the ground truth question, using the question-type taxonomy of Li & Roth (2002). We use GPT-4o (Hurst et al., 2024) to categorize the questions into the main taxonomy categories: entities, humans, numeric, and locations, and the subcategories of ‘descriptions’: manner, reasoning, definition, and description due to their high frequency in the data. We measure agreement in category assignments using pairwise Cohen’s Kappa. Section˜D.1 includes further details on the classification process and on the agreement of GPT-4o with manual human annotations.
BLEU The BLEU surface similarity score (Papineni et al., 2002) between the generated question and the true question.
BertScore (Zhang et al., 2020) provides a cosine similarity score between the generated question and the true question based on contextual embeddings extracted from RoBERTa (Liu et al., 2019).
QA accuracy We introduce a downstream evaluation method with the following logic: the closer the generated question is to the true question, the easier it should be to select the correct answer for the true question when replaced with the generated question. We consider a question as a valid reconstruction of the true question if and only if, given this question, the text, and the four answers for the true question, the model selects the correct answer. We implement this evaluation using RoBERTa (Liu et al., 2019) fine-tuned for multiple choice QA on RACE (Lai et al., 2017), on the 86% of OneStopQA questions that RoBERTa answers correctly. We chose this model because OneStopQA is not in its training data, thus avoiding possible data contamination with newer models.
Baselines
To facilitate model performance interpretation, we introduce the following baseline questions:
Incorrect human questions The two additional human-composed questions for each text available in OneStop, one for a different critical span and one for the same critical span as the true question.
Arbitrary question from an LLM We generate an arbitrary question for the text with GPT-4o, using the prompt in Section˜D.2.
Question from a text-only question decoding model An additional question is generated by a Llama 3.1 model trained to decode the correct question without eye movement data.
6 Results
6.1 Question Selection
Table˜2 presents the aggregated test accuracy across the 10 data splits for the three-way ‘All’ selection, as well as a breakdown into questions with Different Spans, and questions with the Same Span evaluations. A division of the results into the New Text, New Participant, and New Text & Participant regimes is reported in Tables˜4, 5 and 6 in Appendix˜E. Validation set results are reported in Tables˜7, 8, 9 and 10 in Appendix˜E.
Table 2: Question selection test accuracy aggregated across 10 cross-validation splits, with 95% confidence intervals. A majority question choice is identical to chance because the data is balanced across questions. As the data is not i.i.d. (each participant reads multiple paragraphs, and each paragraph is read by multiple participants), we test for differences between models using a linear mixed effects model with random intercepts and slopes for participants and paragraphs: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. The accuracy of all the models is compared to chance performance, where significant gains over this baseline are marked with ‘*’ $p<0.05$, ‘**’ $p<0.01$, and ‘***’ $p<0.001$. The best performing model in each evaluation regime is marked in bold. Significant drops compared to the best model are marked with ‘+’ $p<0.05$, ‘++’ $p<0.01$ and ‘+++’ $p<0.001$.
Model Type Model All
Different Spans
Same Span
Discriminative Chance / Majority Baseline $33.0\pm 0.4_{\text{+++}}$ $55.3\pm 0.4_{\text{+++}}$ $49.9\pm 0.4_{\text{+++}}$
@@ 6.1 Question Selection
Haller RNN (Haller et al., 2022) $41.8\pm 0.6^{***}_{\text{+++}}$ $65.6\pm 0.5^{***}_{\text{+++}}$ $52.1\pm 0.5^{**}_{\text{+++}}$
RoBERTEye-Fixations (Shubi et al., 2024) $\mathbf{49.3}\pm 0.3^{***}$ $\mathbf{70.9}\pm 0.5^{***}$ $\mathbf{57.3}\pm 0.5^{***}$
@@
DalEye-Llama $39.2\pm 0.9^{***}_{\text{+++}}$ $61.4\pm 0.9^{***}_{\text{+++}}$ $52.5\pm 0.4^{***}_{\text{+++}}$
Both Reading-Time Informed Embedding Similarity baseline models perform at chance level across all three evaluation regimes. A possible reason for this outcome lies in the inherent noise in the distribution of reading times of a single participant over a single paragraph, which deems heuristic methods that rely only on this distribution to be ineffective. Haller RNN and RoBERTEye-Fixations perform well above chance, and outperform the generative models on all the evaluations. In the generative models category, DalEye-Llama outperforms DalEye-LLaVA, where the latter is at chance level. Table˜11 in Section˜E.1.1 suggests that the alternative prompt and eye movements representations for DalEye-Llama weaken its performance. The advantage of the discriminative models over the generative models is not surprising, as differently from the discriminative models, the generative models are not explicitly trained on the question selection task.
@@
RoBERTEye-Fixations achieves the highest accuracy across all three evaluation regimes. In Table˜4 of Appendix˜E, we observe that this performance advantage is consistent across the New Text, New Participant, and New Text & Participant evaluations. Thus, the model is able to generalize not only to new participants but importantly also to new texts. Notably, it is also the only model that is well above chance in the Same Span regime with $57.3\%$ accuracy. On the one hand, this confirms that distinguishing between questions that ask about the same portion of the paragraph is indeed much more challenging than distinguishing between questions over different parts of the paragraph. At the same time, the above change performance of RoBERTEye-Fixations on the Same Span evaluation speaks to the fine-grained nature of the information that can be extracted about the reader’s information seeking goals from their eye movements alone.
@@ 6.2 Question Reconstruction
Based on the question selection results in Table˜2, we focus on the DalEye-Llama model in the evaluation of question reconstruction quality. Figure˜4 presents the evaluations in the New Participant and the New Text regimes. Table˜12 in Appendix˜E presents these results in numerical form, and additionally includes the New Text & Participant regime, which yields very similar results to the New Text regime. Figure˜7 in appendix˜E presents an example of DalEye-Llama outputs for the reconstruction task. Table˜13 in Appendix˜E presents reconstruction results for the alternative task prompt for DalEye-Llama, which yields similar results, and the alternative eye movement representations, which yield inferior results to those presented below.
Figure 4: Question reconstruction evaluations of DalEye-Llama for (1) the identity of the generated question word, (2) UIUC semantic category of the question, (3) BLEU score, (4) BERTScore, (5) downstream QA accuracy based on the answer selection of a multiple-choice QA model. DalEye-Llama performance is benchmarked against two human-composed questions for a different and the same critical span, an arbitrary question generated with GPT-4o, and a question from a text-only Llama 3.1 model trained on the question decoding task using only the text. Presented are means with 95% confidence intervals (bootstrapped, $n=1000$).
Importantly, in the New Participant evaluation, DalEye-Llama generated questions outperform all the baseline questions on all five metrics. This suggests that DalEye-Llama generalizes relatively well to new participants when the text appeared in the training set. However, its performance drops both in absolute terms and relative to the baselines in the more challenging regime involving new texts. Notably, in this regime the incorrect human question - same critical span has the highest BLEU, BERTScore, and QA accuracy. This is likely the result of lexical overlap stemming from sharing the critical span with the true question. However, this span-sharing constraint may hinder this baseline in the Question word and Question category evaluations, as sharing a span tends to yield questions that differ in these regards. The human question over a different span, which tends to perform best in these evaluations, is not subject to this constraint.
@@ 7 Related Work
We build on two primary lines of work: (i) psycholinguistic investigation of goal driven reading; and (ii) predictive models of cognitive state from eye movements in reading.
7.1 Analyses of Goal Based Reading
@@
While most prior work in the psychology of reading and psycholinguistics focuses on ordinary reading for general comprehension Rayner et al. (2012), goal or task based reading has been acknowledged as a key frontier in this area (Radach & Kennedy, 2004). Several studies found differences in eye movements between ordinary reading and tasks that define a reading manner: skimming, speed reading and proofreading (Just et al., 1982; Kaakinen & Hyönä, 2010; Schotter et al., 2014; Strukelj & Niehorster, 2018; Chen et al., 2023). Additional studies have found differences between ordinary reading and tasks of target word search (Rayner & Raney, 1996; Rayner & Fischer, 1996), word verification (Radach et al., 2008), and topic search (White et al., 2015). Prior work also analyzed eye movements during human linguistic tasks, such as named entity annotation (Tomanek et al., 2010; Tokunaga et al., 2017). Differences in reading patterns were also found when readers adopt different perspectives on a given text (Kaakinen et al., 2002; Kaakinen & Hyönä, 2008), or given different learning goals (Rothkopf & Billington, 1979). Overall, these works consistently found a systematic influence of the reading task on the resulting reading patterns.
@@
Our work builds most directly on Kaakinen et al. (2015), Malmaud et al. (2020), Hahn & Keller (2023), and Shubi & Berzak (2023), who compared eye movements in ordinary reading and open-ended information seeking. These studies showed task conditioning effects on reading patterns, especially as a function of the task-relevance of the textual information. Similarly to these studies, we formulate the information seeking task as a reading comprehension question specific to the text. We further leverage the previously observed task conditioning effects for automatic discrimination between different tasks over the same text.
7.2 Decoding Cognitive State from Eye Movements
Multimodal modeling of eye movements and text has been a growing area of research in recent years (Reich et al., 2025), with applications ranging from improving NLP (e.g. Klerke et al., 2016; Mishra et al., 2016; Barrett & Hollenstein, 2020; Sood et al., 2020; Deng et al., 2024; López-Cardona et al., 2025) to prediction of eye movement trajectories (e.g. Hollenstein et al., 2021a; Bolliger et al., 2023; Deng et al., 2023). Within this broad line of research, several studies focused on cognitive state decoding: prediction of language-related aspects of the reader and their cognitive state from their eye movements while reading. These include studies on prediction of the reader’s linguistic background (Berzak et al., 2017; Reich et al., 2022; Skerath et al., 2023), language proficiency (Berzak et al., 2018), reading comprehension (Ahn et al., 2020; Reich et al., 2022; Shubi et al., 2024), and subjective readability (Reich et al., 2022).